# The Early Bird Catches the Leak: Unveiling Timing Side Channels in LLM Serving Systems

### v4

@@ IV-B Analysis on PNA
Representative requests are those that most closely approximate the rest of the request space. To identify them, we use the distilbert-base-uncased [67] model by default to generate embeddings and then compute the L2 distance between these embeddings. We then sort the requests by their L2 distances; those with the smallest distances are deemed the most representative and selected as attack requests to maximize coverage. To further expand coverage, we incorporate orthogonal requests—requests that are semantically distinct from one another. This reduces the chance that semantic overlaps among the attacker’s own requests degrade accuracy in identifying victim requests. We classify a cache access as a hit if at least one of the attacker’s requests triggers a cache hit in the timing channel. Although this strategy boosts coverage, it also raises the FPR, necessitating a careful balance.
Figure 10 depicts our greedy search algorithm for locating the most representative requests within a semantically similar target space, thereby improving PNA accuracy. Specifically, during each iteration, we pick the most representative candidate (e.g., $\alpha$) and add it to the attacker’s requests unless it is overly similar to existing ones (e.g., $\beta$). This process continues until no additional candidates remain or until the FPR exceeds a predefined threshold $\sigma$.
In the evaluation, each of the 4 victim request types was tested 500 times. In each iteration, a random pair of private attributes in the Type-1 request was selected. We set $\sigma=0.06$ and identified 5 orthogonal attack requests using the proposed greedy search strategy. Table IV summarizes the TPR for true samples, and the separate FPRs for each false sample type when increasing the number of attack requests from 1 to 5. Specifically, we successfully recovered 407 victim requests out of the 500 true samples with a single attack request, achieving a recovery accuracy of 81.4% with an average FPR of 0.045. With 5 attack requests, 477 victim requests are recovered, demonstrating a recovery accuracy of 95.4% with an average FPR of 0.056. We provide a demo for this attack in our website [62].
TABLE IV: Attack accuracy for the 4 types of victim requests with different number of attack trails.
@@ IV-C Inferring Documents on Commodity LLM
The KV cache can be shared among the same user, within an organization, or even across organizations [31], creating the potential for cross-user information leaks. In our research, we discovered that such leaks are feasible even in remote attack scenarios, particularly when the target LLM processes documents. To demonstrate this, we utilized a document summarization application powered by a commodity LLM API service, where all user requests are processed using the same developer’s API key, with both the victim and the attacker are standard users of the application. We provide an end-to-end demonstration showing how an adversary could infer processed documents from the application through the KV cache side channels. Importantly, such cross-user attacks fundamentally do not require application-specific vulnerabilities, as the KV cache can be shared across organizations [31]. However, our use of the application highlights the privacy risks inherent to LLM-powered applications, even when they fully comply with the guidelines provided by the LLM service.
Note that the observation of the document uploaded to an LLM service, even when the content of the document is known, can expose an organization’s interest, with substantial privacy and competitive ramifications across various domains. For example, a law firm relying on an LLM-based document processor could unknowingly disclose its involvement in complex litigation or pivotal mergers, tipping off opposing parties about strategic decisions before they become public. Similarly, an investment firm analyzing financial statements might inadvertently signal which companies it views as high-potential opportunities, allowing competitors to anticipate emerging deals or investment moves.
The victim application. We implemented the document summarization application with direct summarization [68] (also known as the stuff approach [69]), using the public Deepseek-chat model as the backend LLM API server. The application was built in accordance with Deepseek’s guideline and operates by first extracting text from user-uploaded PDF files using the pdfplumber package. This text is then formatted into a request and sent to the LLM for summarization. In this setup, documents uploaded by users are included in messages under the “user” role and sent to the Deepseek model, which returns summaries of the documents. Notably, all user inputs are processed under a single Deepseek account, which is typical in such applications. However, this poses a privacy concern because Deepseek’s prompt caching [70] can inadvertently allow cached content to be reused across different users. As a result, a malicious user could potentially infer which documents other users have processed by exploiting timing side channels.
@@
Specifically, the victim first submits 100 documents from the interested set. The attacker then probes each of the 200 interested documents once, recording response latencies. Based on a predefined threshold (2.0 seconds in our experiments), the TPR is computed as the fraction of probed documents correctly identified as cache hits (i.e., with latencies below the threshold). Next, the victim uploads 100 additional documents outside the interested set, and the attacker probes the entire set of 200 documents again. The FPR is then calculated as the fraction of documents incorrectly labeled as hits. Our evaluation shows that this attack achieves an average accuracy of 89%, with an average FPR of 0.05. A demo for the attack is presented in our website [62].
Notes. For ethical reasons, we did not explore techniques for forcibly evicting cache entries in real-world systems. Such research would require extensive experimentation, potentially violating usage policies or interfering with other users’ experience. Without active cache eviction, the timing-based attack primarily operates at the granularity where caches naturally expire due to inactivity—around 5 minutes for systems like OpenAI and Anthropic, as indicated in their documentation [70, 24].
@@ IV-D Measurement Study on Commodity LLMs
Specifically, we first measured the response latencies by sending initial requests that were likely to miss the cache. Then, we sent identical requests multiple times and measured the average latencies for these subsequent requests. To maximize the likelihood of co-locating on the same physical machine and ensuring the requests were cached, we conducted continuous tests within the same time period. If we observed lower latencies in the later requests, this indicated the use of caching mechanisms in the LLM services. With KV cache sharing, the computation of matched prefix tokens during the prefill phase can be ignored. However, the output generated during the decoding phase still requires computation and inference, which are influenced by parameters such as temperature, introducing randomness. To verify that the latency reduction was due to KV cache sharing, we deliberately set a high temperature ($0.9$) in the request. This configuration was critical because semantic caching mechanisms typically return identical cached outputs for semantically similar inputs. By introducing substantial randomness in token selection through high temperature, we ensured the LLM would generate diverse responses despite input similarities. We verified whether the LLM produced different responses for each request, with TTFT reductions consistently observed. If it did, this strongly indicated that KV cache sharing was supported, enabling a reduction in TTFT while still allowing for diverse outputs. To minimize the impact of network latency, we sent requests of varying lengths, ranging from 200 to 2,000 tokens. The time difference between cached and uncached responses typically spanned several hundred milliseconds, making it easy to distinguish between the two. Additionally, we observed when the cache is hit, the TTFT remains consistent, regardless of the request length, whereas when the cache is missed, TTFT increases almost linearly as the length of the request grows. As summarized in Table V, most popular LLM service providers support KV cache sharing in specific scenarios.
[b] LLM service System prompt sharing User prompt sharing GPT-4o-mini $\dagger$ Deepinfra Deepseek-chat Claude-3.5 Qwen-max Moonshot Baidu Ernie-8k Google Gemini Fireworks.ai Groq [72] SiliconFLow
TABLE V: Summary of KV cache sharing in real world LLM serving systems (date: 08/29/2024).
  • $\dagger$
    We observed a timing difference on 08/29/2024 and reported it to OpenAI. By late December 2024, the timing difference was no longer stable, despite the API indicating that the prompt cache was effective.
TABLE VI: Native support of semantic caching of popular AI service providers (date: 08/29/2024).
@@
Azure OpenAI Service models [73]
Amazon Bedrock [74]
Google Vertex AI [75]
Alibaba Elastic Algorithm Service (EAS) of Platform for AI (PAI) [76]
Semantic cache sharing. We manually reviewed the documentation of public cloud AI service providers to verify whether they support semantic cache APIs. As shown in Table VI, semantic caching is supported by major AI platform-as-a-service providers. Notably, even on platforms that do not offer native semantic caching, users can still implement their own solutions or leverage open-source alternatives, such as GPTCache.
@@ V-A Mitigating KV Cache Leakages
Design. A straightforward approach to mitigate KV cache leakages is to eliminate any sharing across requests. However, this would negate the computational and cost savings associated with caching. Noting that PSA recovers the request token by token by observing per-token timing differences, we explore the effect of a simple mitigation strategy that the prefix cache can only be shared in units of at least $K$ tokens ($K=2,3,4$ etc.). In SGLang, this could be achieved by modifying the radix tree structure used to manage the KV cache. To prevent the leakages of cache-aware task scheduling, the scheduling policy in SGLang needs to be modified to prioritize requests that have at least $K$ shared prefix tokens. Requests with fewer than $K$ shared prefix tokens would still be placed in the waiting list. This approach reduces the likelihood of KV cache sharing, but it is unlikely to significantly impact performance.
Evaluation. To evaluate the effectiveness of the mitigation and reduce the cost of querying the LLM, we conducted a simulation-based experiment. First, we built the classifiers that detect the hits and misses of $K$ tokens for each value of $K$ ($K=1,2,3,4$), following the method outlined in Section IV-A. Since the timing differences become more pronounced as $K$ increases, we reduced the number of samples (i.e., $n$) in multi-sampling, and obtained the corresponding TPRs and FPRs for each classifier. Then we used an oracle to simulate the classifiers by randomly sampling a number and determining whether it fell within the classification range. In this evaluation, we used the same repetitive trails method, dataset and fine-tuned model as described in Section IV-A. The next-token predictor was modified to predict the next $K$ tokens. Table VII presents the token recovery rate and the average number of queries needed to recover 1 token for $K=1,2,3$ and $4$. The results show that as $K$ increases, the attack still achieves a notable recovery rate. For successfully recovered tokens, the average number of queries decreases with larger $K$ values, as the predictor quickly identifies tokens in easier-to-predict prompts. However, the overall recovery rate declines because the predictor increasingly struggles with harder-to-predict tokens at higher K values. When accounting for tokens not recovered within the maximum 80 allowed guesses, the average query cost per token recovery increases significantly.
TABLE VII: Token recovery results under different numbers of minimum shared tokens.
@@ V-B Mitigating Semantic Cache Leakages
Evaluation. In our prototype implementation, we used the Presidio tool by Microsoft [77] to automatically identify private attributes. For performance evaluation, we used the evaluation dataset released by Presidio, which includes sentences containing private information. Specifically, we randomly sampled 1,000 sentences from the dataset and fed them into both the original GPTCache and the enhanced GPTCache. Then we measured the average delay introduced by the pre-processor and post-processor. The results show that the anonymization process adds an average delay of approximately 6 ms, while GPTCache’s response latency for a semantic cache hit without anonymization is around 0.14 s. Thus, de-identification introduces only about 4% additional overhead, which has a minimal impact on GPTCache’s overall performance.
VI Discussions
Co-locations. Co-location is a prerequisite for timing side-channel attacks. There has been significant research on how to achieve co-location in the cloud for micro-architectural side channels [78, 79, 80, 81]. Co-location can be more efficient in LLM systems because these systems often focus on optimizing cache reuse through improved scheduling policies. For example, baseline scheduling policies, such as first-come-first-serve, typically do not consider the shared prompts within an LLM system. As a result, requests may be mixed across different LLM engines, preventing the reuse of common prompt prefixes and potentially leading to suboptimal cache efficiency. To address this, LLM serving systems like SGLang [26], Parrot [82], Mooncake [83] and BatchLLM [84] have introduced modified schedulers that prioritize requests that align with cached prefixes, a strategy known as cache-aware scheduling.
Unexplored timing leakages. This paper utilizes SGLang [26] and GPTCache [28] as the most representative KV cache and semantic cache sharing mechanisms. However, it is important to note that more sophisticated optimization techniques may enhance performance, but they could also amplify the significance of timing leakage. For example, modular prefix caching [85] and CacheBlend [86] facilitate KV caching for not only prefix tokens but also intermediate ones. Cascade inference [87] stores the shared KV cache in GPU shared memory (SMEM for short), for fast access in multiple requests. The sharing of KV cache in speculative decoding [88] may also introduce speculative side channels, akin to Spectre [89]. We leave the further exploration of the impact of the discovered side channels to future work.
Real-world attacks. In real world attacks, especially targeting online LLM services, the network latency will bring additional noises to the timing. The noise effect could be reduced with more attack trials. Besides, it has been shown possible to mount cache attacks from the remote network [41, 42].
VII Related Works
@@
Prompt extraction attacks with adversarial prompts. Most existing research focuses on stealing system prompts from LLMs by eliciting the previous prompts, typically through direct output or translation. For example, a twitter user claimed to have discovered the prompt used by Bing Chat [90]. Earlier studies involved manually constructing attacker prompts [91, 92], while more recent work, such as PLeak, leverages output feedback from a given prompt and introduces an incremental search algorithm to optimize prompt retrieval [93]. Zhang et al. present a framework for systematically evaluating the effectiveness of these attacks. While these approaches exploit the model’s vulnerability to adversarial prompts, our proposed attacks take advantage of timing differences introduced in LLM service deployment. As such, our attacks do not rely on the specific details of any particular LLM.
@@
Side channel attacks on LLM. Debenedetti et al. propose system-level side channels within the deep learning lifecycle, such as training data filtering, input preprocessing, output monitoring, and query filtering. These side channels can potentially be exploited to infer the training data or the requests [94]. LLM keystroking attacks [95] are a type of packet analysis based side channels. These attacks exploit the length of response tokens, assuming the attacker has access to encrypted network packets. Besides, Carlini et al. investigated the privacy risks of timing attacks in multi-turn interactions with LLM chatbots [96]. By comparison, we are the first to study the timing leaks introduced by LLM serving system optimizations, rather than relying on output data or token packet sizes to recover requests.
@@
VIII Conclusions
LLM inference is a resource-intensive process, prompting numerous studies focused on reducing inference costs and latency. These optimizations often involve the use of various caches. When multiple users share the LLM system, these optimizations can lead to interference between users. This paper examines the side channels created by such interference, identifying two types of leaks: one in the KV cache and another in the semantic cache. We urge LLM system providers to recognize this emerging threat and prioritize security in their design choices.

### v5

@@ IV-B Analysis on PNA
Representative requests are those that most closely approximate the rest of the request space. To identify them, we use the distilbert-base-uncased model by default to generate embeddings and then compute the L2 distance between these embeddings. We then sort the requests by their L2 distances; those with the smallest distances are deemed the most representative and selected as attack requests to maximize coverage. To further expand coverage, we incorporate orthogonal requests—requests that are semantically distinct from one another. This reduces the chance that semantic overlaps among the attacker’s own requests degrade accuracy in identifying victim requests. We classify a cache access as a hit if at least one of the attacker’s requests triggers a cache hit in the timing channel. Although this strategy boosts coverage, it also raises the FPR, necessitating a careful balance.
Figure 10 depicts our greedy search algorithm, which iteratively selects the most representative candidates (e.g., $\alpha$) within a semantically similar target space, unless they are overly similar to previously selected ones (e.g., $\beta$). This process continues until no suitable candidates remain or when the FPR exceeds a predefined threshold $\sigma$.
In the evaluation, each of the 4 victim request types was tested 500 times. In each iteration, a random pair of private attributes in the Type-1 request was selected. We set $\sigma=0.06$ and identified 5 orthogonal attack requests using the proposed greedy search strategy. Table III summarizes the TPR for true samples, and the separate FPRs for each false sample type when increasing the attack requests from 1 to 5. Specifically, we recovered 407 victim requests out of the 500 true samples with only 1 attack request, achieving a recovery accuracy of 81.4% with an average FPR of 0.045. With 5 attack requests, 477 victim requests were recovered, demonstrating a recovery accuracy of 95.4% with an average FPR of 0.056. A demo of this attack is available on our website [35].
TABLE III: Attack accuracy for the 4 types of victim requests with different number of attack trials.
@@ IV-C Inferring Documents on Commodity LLM
The KV cache can be shared among the same user, within an organization, or even across organizations [17], creating the potential for cross-user information leaks. In our research, we discovered that such leaks are feasible even in remote attack scenarios, particularly when the target LLM processes documents. To demonstrate this, we utilized a document summarization application powered by a commodity LLM API service, where all user requests are processed using the same developer’s API key, and both the victim and the attacker are standard users of the application. We provide an end-to-end demonstration showing how an adversary could infer processed documents from the application through the KV cache side channels. Importantly, such cross-user attacks fundamentally do not require application-specific vulnerabilities, as the KV cache can be shared across organizations [17]. However, our use of the application highlights the privacy risks inherent to LLM-powered applications, even when they fully comply with the guidelines provided by the LLM service.
Note that the observation of the document uploaded to an LLM service, even when the content of the document is known, can expose an organization’s interest, with substantial privacy and competitive ramifications across various domains. For example, a law firm relying on an LLM-based document processor could unknowingly disclose its involvement in complex litigation or pivotal mergers, tipping off opposing parties about strategic decisions before they become public.
The victim application. We implemented the document summarization application with direct summarization (also known as the stuff approach [38]), using the public Deepseek-chat model as the backend LLM API server. The application was built in accordance with Deepseek’s guideline and operates by first extracting text from user-uploaded PDF files using the pdfplumber package. This text is then formatted into a request and sent to the LLM for summarization. In this setup, documents uploaded by users are included in messages under the “user” role and sent to the Deepseek model, which returns summaries of the documents. Notably, all user inputs are processed under a single Deepseek account, which is typical in enterprise LLM deployments. However, this poses a privacy concern because Deepseek’s prompt caching can inadvertently allow cached content to be reused across different users. As a result, a malicious user could potentially infer which documents other users have processed by exploiting timing side channels.
@@
Specifically, the victim first submits 100 documents from the interested set. The attacker then probes each of the 200 interested documents once, recording response latencies. Based on a predefined threshold (2.0 seconds in our experiments), the TPR is computed as the fraction of probed documents correctly identified as cache hits (i.e., with latencies below the threshold). Next, the victim uploads 100 additional documents outside the interested set, and the attacker probes the entire set of 200 documents again. The FPR is then calculated as the fraction of documents incorrectly labeled as hits. Our evaluation shows that this attack achieves an average accuracy of 89%, with an average FPR of 0.05. A demo for the attack is presented on our website [35].
Discussions. Since KV-cache sharing typically requires requests to have a common prefix, we assume that the attacker knows either the full document or at least its prefix (i.e., the initial portion of the content). However, recent work has introduced more flexible cache-sharing mechanisms that extend beyond strict prefix matching. For example, CacheBlend [40] and LMCache [41] enable sharing of arbitrary reused text segments, while KVShare [42] and SemShareKV [43] support semantic-aware key-value cache sharing. These techniques are already being integrated into popular LLM serving engines such as vLLM, llm-d, and KServe. In such systems, an attacker could probe whether a target document has been processed by the LLM using only partial common text (not necessarily a prefix) or even general topical knowledge. We leave deeper exploration of these attack scenarios to future work.
@@ IV-D Measurement Study on Commodity LLMs
Specifically, we first measured the response latencies by sending initial requests that were likely to miss the cache. Then, we sent identical requests 10 times and measured the median latencies for these subsequent requests. To maximize the likelihood of co-locating on the same physical machine and ensuring the requests were cached, we conducted continuous tests within the same time period. If we observed lower latencies in the later requests, this indicated the use of caching mechanisms in the LLM services. With KV cache sharing, the computation of matched prefix tokens during the prefill phase can be ignored. However, the output generated during the decoding phase still requires computation and inference, which are influenced by parameters such as temperature, introducing randomness. To verify that the latency reduction was due to KV cache sharing, we deliberately set a high temperature ($0.9$) in the request. This configuration was critical because semantic caching mechanisms typically return identical cached outputs for semantically similar inputs. By introducing substantial randomness in token selection through high temperature, we ensured the LLM would generate diverse responses despite input similarities. We verified whether the LLM produced different responses for each request, with TTFT reductions consistently observed. If it did, this strongly indicated that KV cache sharing was supported, enabling a reduction in TTFT while still allowing for diverse outputs. To minimize the impact of network latency, we sent requests of varying lengths, ranging from 200 to 2,000 tokens. The time difference between cached and uncached responses typically spanned several hundred milliseconds, making it easy to distinguish between the two. Additionally, we observed when the cache is hit, the TTFT remains consistent, regardless of the request length, whereas when the cache is missed, TTFT increases almost linearly as the length of the request grows. As summarized in Table IV, most popular LLM service providers support KV cache sharing in specific scenarios.
[b] LLM service System prompt sharing User prompt sharing GPT-4o-mini Deepinfra Deepseek-chat Claude-3.5 Qwen-max Moonshot Baidu Ernie-8k Google Gemini Fireworks.ai Groq SiliconFLow
TABLE IV: Summary of KV cache sharing in real world LLM serving systems (date: 08/29/2024).
TABLE V: Native support of semantic caching of popular AI service providers (date: 08/29/2024).
@@
Azure OpenAI Service models
Amazon Bedrock
Google Vertex AI
Alibaba Elastic Algorithm Service (EAS) of Platform for AI (PAI)
Semantic cache sharing. We manually reviewed the documentation of public AI service providers to verify whether they support semantic cache APIs. As shown in Table V, semantic caching is supported by major AI platform-as-a-service providers. Even on platforms that do not offer native semantic caching, users can still implement their own solutions or leverage open-source alternatives, such as GPTCache.
@@ V-A Mitigating KV Cache Leakages
Design. A straightforward approach to mitigate KV cache leakages is to eliminate any sharing across requests. However, this would negate the computational and cost savings associated with caching. Noting that PSA recovers the request token by token by observing per-token timing differences, we explore the effect of a simple mitigation strategy that the prefix cache can only be shared in units of at least $K$ tokens ($K=2,3,4$ etc.). In SGLang, this could be achieved by modifying the radix tree structure used to manage the KV cache. This approach reduces the likelihood of KV cache sharing, but it is unlikely to significantly impact performance.
Evaluation. To evaluate the effectiveness of the mitigation and reduce the cost of querying the LLM, we conducted a simulation-based experiment. First, we built the classifiers that detect the hits and misses of $K$ tokens for each value of $K$ ($K=1,2,3,4$), following the method outlined in Section IV-A. Since the timing differences become more pronounced as $K$ increases, we reduced the number of samples (i.e., $n$) in multi-sampling, and obtained the corresponding TPRs and FPRs for each classifier. Then we used an oracle to simulate the classifiers by randomly sampling a number and determining whether it fell within the classification range. In this evaluation, we used the same repetitive trials method, dataset and fine-tuned model as described in Section IV-A. The next-token predictor was modified to predict the next $K$ tokens. Table VI presents the token recovery rate and the average number of queries per recovered token for $K=1,2,3$ and $4$. The results show that increasing $K$ still yields a notable recovery rate, but the attack bottleneck shifts fundamentally. Larger $K$ amplifies timing differences between cache hits and misses, reducing false positives. At the same time, the predictor’s search space grows exponentially, making predictor accuracy the new bottleneck. For tokens that are successfully recovered, the average number of queries decreases with larger $K$ values, since the predictor can more easily identify tokens in simpler prompts. However, the overall recovery rates decline because the predictor struggles with harder tokens at higher $K$. As a result, although queries per recovered token drop, the total cost of recovering an entire prompt rises significantly.
TABLE VI: Token recovery results under different numbers of minimum shared tokens.
@@ V-B Mitigating Semantic Cache Leakages
Evaluation. In our prototype implementation, we used the Presidio tool [44] to automatically identify private attributes. For performance evaluation, we used the evaluation dataset released by Presidio, which includes sentences containing private information. Specifically, we randomly sampled 1,000 sentences from the dataset and fed them into both the original GPTCache and the enhanced GPTCache. Then we measured the average delay introduced by the pre-processor and post-processor. The results show that the anonymization process adds an average delay of approximately 6 ms, while GPTCache’s response latency for a semantic cache hit without anonymization is around 0.14 s. Thus, de-identification introduces only about 4% additional overhead, which has a minimal impact on GPTCache’s overall performance.
VI Related Works
@@
Prompt extraction attacks with adversarial prompts. Most existing research focuses on stealing system prompts from LLMs by eliciting the previous prompts, typically through direct output or translation. Earlier studies involved manually constructing attacker prompts [45, 46], while more recent work, such as PLeak, leverages output feedback from a given prompt and introduces an incremental search algorithm to optimize prompt retrieval [47]. While these approaches exploit the model’s vulnerability to adversarial prompts, our proposed attacks take advantage of timing differences introduced in LLM service deployment. As such, our attacks do not rely on the specific details of any particular LLM.
@@
Side channel attacks on LLM. Debenedetti et al. propose system-level side channels within the deep learning lifecycle, such as training data filtering, input preprocessing, output monitoring, and query filtering. These side channels can potentially be exploited to infer the training data or the requests [48]. LLM keystroking attacks [49] are a type of packet analysis based side channels. These attacks exploit the length of response tokens, assuming the attacker has access to encrypted network packets. Besides, Carlini et al. investigated the privacy risks of timing attacks in multi-turn interactions with LLM chatbots [50]. More recently, Soleimani et al. exposed network side-channel vulnerabilities in speculative decoding-optimized LLM services [51]. By comparison, we are the first to study the timing leaks introduced by LLM serving system optimizations, rather than relying on output data or token packet sizes to recover requests.
@@
VII Conclusions
LLM inference is a resource-intensive task, prompting studies focused on reducing inference latency. These optimizations often involve the use of various caches. When multiple users share the LLM system, these optimizations can lead to interference between users. This paper examines the side channels created by such interference, identifying two types of leaks: one in the KV cache and another in the semantic cache. We urge LLM system providers to recognize this emerging threat and prioritize security in their design choices.